OpenStudy (anonymous):

Divide & Simplify b^4/b^-9

OpenStudy (qqstory):

B^13

OpenStudy (anonymous):

How did you get that answer

OpenStudy (qqstory):

Hacks!

OpenStudy (qqstory):

divide means to subtract in expoential simplication..

OpenStudy (zepp):

Remember the rule of exponents, a^b / a^c Same term, it would be a^(b-c)

OpenStudy (anonymous):

\[\frac{b^m}{b^n}=b^{m-n}\] and so \[\frac{b^4}{b^{-9}}=b^{4-(-9)}=b^{4+9}=b^{13}\]

OpenStudy (qqstory):

YES YES!

OpenStudy (anonymous):

thank you everyone

OpenStudy (zepp):

@qqstory, giving out an answer without explanation is a sin on OpenStudy.

OpenStudy (anonymous):

it helps alot when someone explains it so then i understand it .....but i understand what everyone is saying